### 04 Mar 2024 v1

# A Generative Approach for Wikipedia-Scale Visual Entity Recognition

Abstract
Code will be released.
In this paper, we address web-scale visual entity recognition, specifically the task of mapping a given query image to one of the 6 million existing entities in Wikipedia. One way of approaching a problem of such scale is using dual-encoder models (e.g. CLIP), where all the entity names and query images are embedded into a unified space, paving the way for an approximate $k$NN search. Alternatively, it is also possible to re-purpose a captioning model to directly generate the entity names for a given image. In contrast, we introduce a novel Generative Entity Recognition (ger) framework, which given an input image learns to auto-regressively decode a semantic and discriminative “code” identifying the target entity. Our experiments demonstrate the efficacy of this ger paradigm, showcasing state-of-the-art performance on the challenging OVEN benchmark. ger surpasses strong captioning, dual-encoder, visual matching and hierarchical classification baselines, affirming its advantage in tackling the complexities of web-scale recognition.
1 Introduction
Generative vision-language models such as GPT-4 [30], Flamingo [2] or PALI [5], are becoming increasingly popular for computer vision applications. They show an impressive ability to generate free-form text for describing the contents of an image (captioning), or answering questions based on an image (visual-question answering). Nevertheless, their potential for recognition tasks [12], which usually require a more concise, structured output, remains under-explored. The focus of this paper is to explore their application for the challenging task of web-scale entity recognition. A recent benchmark, Open-domain Visual Entity recognitioN (OVEN) [12], challenges models to associate an image with a Wikipedia entity from a pool of over six million entities. Models must establish a robust association between images across millions of coarse-grained and fine-grained entities, encompassing a wide spectrum of concepts such as animals, buildings, locations, and a multitude of others [12].
Figure 1: We introduce ger, a novel generative paradigm for web-scale visual entity recognition. We create compact semantic codes for each entity, and learn to auto-regressively generate them for a given query image at inference.
Traditionally, the predominant methods employed to address the challenge of visual entity recognition have revolved around either classification or contrastive dual-encoder paradigm like CLIP [32]. While classification offers a straightforward approach, it grapples with limitations when confronted with extensive label spaces such as that of OVEN, resulting in substantial parameter counts and practical engineering complexities. The dual-encoder approach on the other hand, learns a unified image-text feature space, thereby facilitating efficient nearest neighbor searches for recognition. Nonetheless, this approach exhibits its own drawbacks: (a) it does not directly optimize for the final recognition task but instead relies on indirect optimization through contrastive loss where a set of negative data has to be subsampled at training time [11, 29, 32], (b) compressing either the image or text into an embedding vector results in loss of information, detrimentally affecting performance for fine-grained recognition [15] and (c) the memory requirements for storing dense representations scale proportionally with the size of the entity set.
These challenges of the dual-encoder paradigm have kindled interest in alternative strategies. Notably, in Natural Language Processing (NLP) domain, recent works challenge the dual-encoder approach and use generative models instead for information retrieval [42, 33, 31, 6, 25, 41]. These works represent each element of the corpus by a compact code of integers, and learn an auto-regressive generative model to decode the target code for a given query. This paradigm promises to overcome some drawbacks of dual-encoders by simplifying the retrieval pipeline such that the training and inference objectives are the same, and directly encoding the corpus within the model’s parameters. Also as an alternative to dual encoders, OVEN paper [12] showcases the feasibility of extending a generative image captioning model [5] for visual entity recognition by matching the generated caption to one of the Wikipedia entity texts [34].
Inspired by these recent explorations, we propose a Generative Entity Recognition (ger) framework (illustrated in Fig. 1) to facilitate end-to-end visual entity recognition by leveraging generative auto-regressive models. Specifically, we represent each Wikipedia entity with a code, i.e. a short sequence of integers. Then, we train models to predict an entity from an input image by auto-regressively generating the code corresponding to the target entity. We find that creating unAmbiguous, Language-based and Discriminative (ald) entity codes results in the best variant of our ger framework, which we denote by ger-ald. In fact, while we observe that unstructured “atomic” codes work well in some scenarios, they fail when training data or model capacity are limited or more importantly, when the entity set reaches the million scale (see Sec. 4.4.1). Plus, they cannot generalize to new entities. In contrast, we find that semantically-structured codes based on language improve upon atomic codes by leveraging generic concepts shared across related entities (see example in Fig. 1 with “Black colobus” and “Black-and-white colobus” sharing common code tokens). A simple way of creating codes based on language is to directly tokenize [20] the entity name, which is akin to image captioning where the entity name is used as a caption [12, 6]. However, we find that such tokenized entity names contain clutter and noisy information, all the more so when the entity name is long (see Sec. 4.4.2). Our ger-ald method improves over this simple captioning baseline by decoding only the most discriminative part of the tokenized entity name, i.e. the part which makes the considered entity name the most different compared to all other entities.
Finally, we also propose an entity-based pre-training to condition the ger models to web-scale entity recognition. Inspired by recent advances in retrieval-based methods [15, 23], we retrieve a subset of images from a large-scale image-text dataset typically used for captioning or contrastive pre-training [5] and re-purpose it by replacing the original text captions with related OVEN entity names. Overall, our experiments demonstrate the efficacy of the proposed ger paradigm: ger-ald outperforms previously published numbers on OVEN benchmark [12] by $+6.7$ top-1 accuracy, while using 42×42$\times$42 × less parameters. In summary, our contributions are as follows:
  • a generative entity recognition framework (ger) to facilitate end-to-end visual entity recognition;
  • an innovative strategy for encoding Wikipedia entities into unambiguous language-based discriminative (ald) codes that are highly effective for ger;
  • an entity-based pre-training process without requiring human intervention;
  • state-of-the-art results in challenging web-scale OVEN entity recognition and on-par performance to traditional classifiers in smaller-scale label-space scenarios.
2 Related work
Visual entity recognition aims to recognize classes, or entities given visual inputs [35]. Granularity of visual entity recognition tasks varies from every-day generic objects [8, 9], to fine-grained domains, such as birds [44], dogs [17], cars [18], food [4], landmarks [47], faces [50] and natural world species [43]. Some challenges for the visual entity recognition tasks include imbalanced training classes following a long-tailed distribution [24], or noisy training labels [22]. Recent work [12] proposes a new, web-scale dataset for open-domain entity recognition. This challenging benchmark contains $6$M entity names derived from Wikipedia page titles, including coarse-grained and fine-grained entities, encompassing a wide spectrum of concepts such as animals, buildings, organizations, landmarks, and a multitude of other. The authors show that generative captioning models (i.e. PaLI [5]) outperform dual encoder models for large-scale entity recognition. In this paper, we build upon this observation, and study generative models for accurate and efficient entity recognition.
Extreme classification tackles entity recognition specifically at a very large scale with a pure classification approach [3, 26, 1]. Typical approaches explore strategies for scaling to the hundred of thousands scale and preliminary results are even shown at million scale [1]. By leveraging generative image-to-text models, we propose a fresh perspective beyond traditional classification methods typically used in the context of large-scale visual entity recognition.
Generative auto-regressive retrieval methods are increasingly popular in NLP [42, 33, 31, 6, 25, 41]. GENRE retrieves Wikipedia entities by generating their names in an autoregressive fashion. Seminal work DSI [42] shows the benefit of learning to decode compact codes (created either randomly or with hierarchical k-means clustering) associated with each document. Neural Corpus Indexer [46] proposes a specific decoding scheme for generative retrieval and show the benefit of query augmentation by automatically generating training queries for documents to be indexed. TIGER [33] studies generative retrieval in the context of recommender systems. Finally, [31] conducts a systematic study of generative retrieval systems when scaled to millions of document passages. Only very few works explore this family of approaches in computer vision domain, and only in very small-scale and uni-modal scenarios [49].
3 Method
Figure 2: Overview of ger-ald method. (a) We utilize a text tokenizer to create compact and semantic codes, which represents each entity with short, but discriminative representations. (b) We learn a generative auto-regressive model, which learns to decode the correct code for given query image and text pair.
Our goal is to explore how to adapt generative auto-regressive models to the task of visual entity recognition (ger). While previous works have shown preliminary signal that it is possible to repurpose autoregressive models for entity recognition by directly decoding entity names [12, 6], we propose a more effective strategy. An overview of our framework is in Fig. 2.
3.1 Problem definition
Web-scale visual entity recognition. The Open-domain Visual Entity recognitioN (OVEN) [12] task consists of mapping input visual queries to one of the $6$M English Wikipedia entities. More specifically, for a given image query $x_{v}$ and text query $x_{t}$, the model needs to recognize the corresponding entity $e$ among the set $\mathcal{E}$ of all possible entities. The purpose of the input text $x_{t}$ is to achieve unambiguous recognition. For example, when several entities are represented in the query image $x_{v}$, the text query indicates which one needs to be recognized. Each entity $e\in\mathcal{E}$ comes with an entity name, denoted by $t_{e}$, which corresponds to the title of the entity Wikipedia page.
Representing each entity with a code. In ger, we represent each entity $e$ by a code denoted by $c^{e}=\{c^{e}_{1},...,c^{e}_{L}\}\in\llbracket 1,V\rrbracket^{L}$ where $L$ is the length of the code and $V$ is the size of the vocabulary of all integer values that each code token $c^{e}_{i}$ can take. This forms up to $V^{L}$ unique codes. Note that vanilla image classification and captioning baselines can both be cast into this code formulation. In fact, with $L=1$ and $V=|\mathcal{E}|$, the codes are equivalent to the labels used in standard multi-class classification. On the other hand, if each code token value in $\llbracket 1,V\rrbracket$ maps to a (sub-)word in a pre-defined vocabulary [20], then the codes simply correspond to standard tokenized text used in captioning models [19, 39, 45]. In the following paragraphs, we detail ger-ald, our most effective strategy for building codes $\mathcal{C}$ to represent all 6M English Wikipedia entities.
3.2 ger-ald: Creating ald codes for ger
We design the code set $\mathcal{C}$ so that it has three properties which we find are important for effective ger models: i) semantically structured thanks to language, ii) discriminative and compact, and iii) unambiguous. Our algorithm to create such unambiguous, language-based and discriminative codes, called ald, is illustrated in Fig. 2 (a), described in pseudo-code in Algorithm 1 and detailed in the following.
[…]
A simple yet effective way of having semantic codes is to tokenize the entity names based on text tokenizers [19, 39, 20, 6]. If each of the sub-words in the entity names are mapped to an integer representing this sub-word, then entities Q358813 and Q521977 naturally share code tokens: those representing the phrase “colobus”. We denote by $\Phi(.)$ an off-the-shelf text tokenizer with a vocabulary of $V_{\Phi}$ sub-words such that $\Phi(t_{e})=\{y^{e}_{1},...,y^{e}_{L_{e}}\}\in\llbracket 1,V_{\Phi}\rrbracket^{L_{e}}$ where $L_{e}$ is the length of the tokenized entity name $\Phi(t_{e})$. In practice we use the same language tokenizer as GIT [45] for $\Phi(.)$ and have a vocabulary size of $V=V_{\Phi}=30522$. We refer to the baseline of using codes $\mathcal{C}$ created by simple tokenization of the entity name as ger-caption (i.e. we treat the entity name as a caption) [6]. We show in the following paragraph how ger-ald codes differ from such ger-caption codes by making them more compact and discriminative.
[…]
Unambiguous codes. Note that several entities might share the same least frequent $(L-1)^{\text{th}}$ tokens. In this case their code are exactly identical up to the $(L-1)^{\text{th}}$ token. We use the last $L^{\text{th}}$ token to ensure that each entity has a unique code: we greedily assign the last code token $c^{e}_{L}$ to the next least frequent word of the tokenized entity name until the code $c_{e}$ is different from all existing codes. If this still fails to create a unique code, we assign $c^{e}_{L}$ to a random token value $v^{\prime}$ so that the resulting code is unique. With code length $L=4$, only $0.5\%$ of the entities use a random token value.
3.3 Training
In this section, we describe the model used to decode entity codes from an input image-text pair. Importantly, we also introduce our entity-based pre-training to condition the generative model to the task of entity recognition.
Auto-regressive generative models. We build upon GIT [45], an auto-regressive image-to-text generative model. The query image-text pair $(x_{v},x_{t})$ is transformed into a set of $d$-dimensional embeddings using a visual encoder for $x_{v}$ and the text tokenizer $\Phi(.)$ for $x_{t}$. The resulting output is represented by $\mathbf{X}_{v}\in\mathbb{R}^{N_{v}\times d}$ (resp. $\mathbf{X}_{t}\in\mathbb{R}^{N_{t}\times d}$) for image (resp. text) tokens. We then input $\mathbf{X}_{v}$ and $\mathbf{X}_{t}$ to a decoder network $g(.)$ whose task is to decode the next code token $c^{e}_{i}$, conditioned on the previous tokens $c^{e}_{j<i}$. Each code token value $v$ in $\llbracket 1,V\rrbracket$ maps to a learnable $d$-dimensional vector $\mathbf{Y}_{v}$ (gathered in the embedding matrix $\mathbf{Y}\in\mathbb{R}^{(V+1)\times d}$ where $\mathbf{Y}_{0}$ corresponds to the “beginning of code” token). We train with a language modeling loss:
$\mathcal{L}^{e}=\frac{1}{L}\sum_{i=1}^{L}\ell(c^{e}_{i},g([\mathbf{X}_{v};\mathbf{X}_{t};\mathbf{Y}_{0};\mathbf{Y}_{c^{e}_{0<j<i}}]))\vspace{-0.3cm}$
where [;][;][ ; ] corresponds to the concatenation operation in the first dimension and $\ell$ is the softmax cross-entropy loss with label-smoothing [27]. We average $\mathcal{L}^{e}$ over a mini-batch and learn the weights of the visual encoder, decoder $g(.)$ and embedding matrix $\mathbf{Y}$ through back-propagation. When decoding, we use beam search to obtain the best predicted entity coded. We find that we do not need to constrain the beam search to existing codes since more than $99\%$ of the top-$1$ predictions are valid codes for converged ger models.
Entity-based pre-training. Common auto-regressive models such as GIT [45] or PaLI [5] are pre-trained for descriptive captioning. As shown in Tab. 5 and Fig 9, they generalize poorly to entity recognition. This is because of the task discrepancy between predicting a descriptive caption and predicting an entity name (see examples in Fig. 9). In order to condition our models better for entity recognition, we propose to collect a significant number of entity-based pretraining images, each associated with a Wikipedia entity instead of a generic caption. However, such an entity-based pretraining dataset does not exist. We create it in an automatic way, without any human supervision.
To do so, we leverage existing large-scale image-caption datasets [37, 38]: unless specified otherwise we use WebLI [5]. For each Wikipedia entity, we retrieve in WebLI the image-caption pairs that best represent this entity and replace their original captions by this entity name [15, 23]. Specifically, we embed the $6$M entity names of OVEN with a semantic text encoder [32] and find the top-$k$ most similar captions in WebLI. We retrieve their corresponding images and replace their original captions by the considered entity name. We ensure that no image is assigned to multiple entities to avoid instability during training. We vary the number of retrieved images $k$ per entity from $2$ to $100$ to produce pre-training datasets of different sizes: from 11M up to 55M images (see Fig. 6). We denote by Entity-WebLI (resp. Entity-LAION) the resulting dataset used for entity-based pretraining, built from WebLI (resp. LAION [38]). This way of creating pre-training data is akin to the query generation techniques used for generative retrieval in NLP [46]. However, rather than generating a synthetic input, we simply retrieve input images from a large-scale dataset.
3.4 Baselines
We compare our method to the following different baselines.
Hierarchical classification. Solving million-scale entity recognition with classification is unpractical due to the very large number of classes. A workaround is to use hierarchical classifiers. As OVEN does not come with hierarchical labels we obtain a 3-level hierarchy through k-means of the 6M entity names encoded with sentence-T5 [28]. We train a multi-class classifier for each parent node in the hierarchy. To avoid training a huge number of different classification matrices, we learn a generic classifier matrix per level which is modified by learnable small modifiers depending on the path in the hierarchy.
Dual encoders. Another typical workaround to classification is to rely on deep metric learning approaches [36] such as Noise Contrastive Estimation [11] and its InfoNCE variant [29] as used in popular dual encoder approaches [32, 16]. Dual encoders learn a unified image-text feature space with separate encoders, thereby facilitating efficient nearest neighbor searches for recognition. We use CLIP-L/14 [32].
Visual matching. We also experiment with pure visual matching baselines. We use off-the-shelf CLIP-L/14 visual encoder and Entity-WebLI (55M) dataset as the memory. We use $k=500$ for nearest neighbor search with majority voting as it obtains the best results on OVEN val set.
Captioning. We compare to Git-Large [45] or PaLI [5] image-to-text auto-regressive captioning models.
ger-baselines: alternative code creation strategies. We compare ger-ald, i.e. the best variant of ger, with several alternatives. First, ger-atomic refers to using atomic, completely unstructured codes, i.e. each code token $c^{e}_{i}$ is randomly drawn from $\llbracket 1,V\rrbracket^{L}$ [42] . Second, we consider two alternatives using semantically structured codes: (i) ger-hkc where we embed all the entity names with a pretrained text encoder before applying hierarchical k-means clustering on the resulting embeddings [42] and (ii) ger-caption where we create a code by tokenizing the entity name with $\Phi(.)$ [12, 6]. More details on these baselines are in Sec. 6.4.
4 Experiments
In this section, we detail our experimental setup, compare our method with state of the art and baselines, and finally present thorough analyses on code creation and pretraining.
4.1 Experimental setting
OVEN dataset consists of 6,063,945 different entities [12]. We evaluate the models on the validation and test splits, by reporting the harmonic mean (HM) of top-1 accuracy scores between “seen” and “unseen” entities. Seen are entities present in the OVEN training set. Unseen entities are a subset of entities among the ones not present in the training set. The models are evaluated on a total of 3192 entities (1721 for seen and 1471 for unseen) for validation and 15888 entities (8355 for seen and 7533 for unseen) for test. We call the entities that the model is evaluated on by “positive” entities (i.e. the union of the 3192 validation and 15888 test entities) and all other entities by “negative” entities.
Pretraining OVEN test
Method #par.(B) dataset #imgs HM seen unseen
Dual encoder approaches
$\text{CLIP}_{\text{ViT-L14}}$ 0.42 OpenAI 400M  5.2 5.6 4.9
$\text{CLIPfusion}_{\text{ViT-L14}}$ 0.88 OpenAI 400M  8.4 33.6 4.8
[…]
Captioning approaches
GiT-Large 0.40 WebLI 100M  7.0 17.6 4.3
PaLI-3B 3 WebLI 1B  9.1 19.1 6.0
PaLI-17B 17 WebLI 1B 16.0 28.3 11.2
Generative entity recognition
[…]
ger-ald (Ours) 0.40 Entity-WebLI 55M 22.7 31.5 17.7
Table 1: Comparison with state-of-the-art approaches on OVEN entity test split. We report the harmonic mean (HM) of the seen and unseen splits (top-1 accuracy) after finetuning on OVEN training set. Numbers are taken from [12] except methods based on GiT-Large which are run by us. We indicate the total number of parameters of each model (“# par.”) in billion and the pretraining dataset details. $\ddagger$: use only publicly available data.
Pretraining and finetuning. Unless specified otherwise, we pretrain our models on the entity-WebLI dataset, which we create considering all 6M entity names as described in Sec. 3.3. After this entity-based pretraining, the models are finetuned on OVEN training set which consists only of the “seen” entities. All implementation details are in Sec. 6 and code will be released in scenic library [7].
Preventing data leakage. We remove pretraining images from Entity-WebLI and Entity-LAION with a cosine similarity (with CLIP-L/14 visual features) above 0.95 with any of the OVEN test or val images. We chose a 0.95 conservative threshold by looking at some examples: similarity 0.95 corresponds to conceptually similar images but clearly not duplicates (see Fig. 8).
4.2 Comparison with the state of the art
In Tab. 1, we compare the performance of ger-ald, our best ger variant, on the OVEN entity benchmark with previously published numbers after finetuning on the OVEN training set. We see that our method outperforms previously proposed approaches by significant margins. Notably, ger-ald improves over the captioning model PALI-17B by $+6.8$ top-1 HM test accuracy (a relative improvement of $43\%$) while using 42×42$\times$42 × less parameters.
4.3 Comparison with baselines
In Tab. 2, we compare ger-ald with the different baselines described in Sec. 3.4. All baselines use exactly the same pretraining dataset entity-based WebLI (55M) and model architectures of comparable sizes.
Comparing ger to different paradigms. We see in Tab. 2 that ger outperforms strong captioning, dual-encoder, visual matching and hierarchical classification baselines, affirming its advantage in tackling web-scale visual entity recognition. Our superior performance compared to dual encoders aligns with previous works observing that CLIP struggles for fine-grained recognition [12, 15]. Due to query image and entity name similarities being captured only through a vector dot product, potentially fine-grained interactions are missed. Also, ger offers significant advantages over dual encoders: its computational complexity is not a function of entity set size and it does not require to store entity dense embeddings.
Different ger variants. In Tab. 2, we compare different variants of ger: one variant using unstructured codes (ger-atomic) and three variants using semantically-structured codes: ger-caption, ger-hkc and ger-ald. We observe that ger-ald is the best performing variant, both after entity-based pretraining and after finetuning on the OVEN seen entities. Compared to ger-caption, ger-ald use codes that are more discriminative and compact, which improves the performance particularly for entities with long names (see Sec. 4.4.2). Compared to ger-atomic, ger-ald codes yield a semantic structure which is crucial for million-scale label-space as shown in Sec. 4.4.1. ger-hkc model also gets strong performance but relies on an off-the-shelf semantic text encoder which makes the approach more complex and costly compared to ger-ald. ger-hkc is a first step towards learning codes and we hope future works will propose original and better code creation strategies [41].
Entity-based pretraining + finetuning on seen
Method HM seen unseen HM seen unseen
Dual encoders  9.2 8.9 9.4 16.3 24.3 12.3
Visual matching 16.2 15.5 17.1 16.4 15.7 17.2
Captioning 13.2 13.1 13.3 16.8 25.9 12.5
Hierarchic. classif. 14.7 14.8 14.6 21.8 29.6 17.2
Generative entity recognition
ger-atomic 15.9 15.3 16.7 20.1 26.2 16.3
ger-caption 14.3 16.5 12.6 20.7 26.8 16.9
ger-hkc 15.8 15.5 16.0 21.0 25.2 17.9
ger-ald 17.7 18.3 17.2 22.7 31.5 17.7
Table 2: Baseline comparisons. All baselines use exactly the same pretraining dataset Entity-WebLI (55M) and architectures of comparable number of parameters ($\sim 400$M). All numbers are obtained with finetuning on seen split after entity-based pretraining. We report the Harmonic Mean of top-1 accuracy on OVEN test.
4.4 Analysis and ablation study
In this section, unless specified otherwise, we report the accuracy on the OVEN validation set [12] evaluated after pretraining on Entity-WebLI (27M), i.e. no OVEN finetuning.
4.4.1 Semantic versus atomic codes
In Fig. 3 (and Tab. 6), we report the relative improvement of semantically-structured codes (ger-ald) compared to unstructured codes (ger-atomic). We vary: (i) pretraining data size, (ii) model capacity and (iii) label-space size. A relative improvement of $100\%$ means that the performance of ger-ald is doubled compared to ger-atomic.
[…]
Limited pretraining data. In Fig. 3 (left), we see that semantic codes outperform atomic codes when the amount of data available for pretraining diminishes. In fact, the results reported in Tab. 2 corresponds to the most favorable scenario for ger-atomic with 55M pretraining datapoints (represented by in Fig. 3). The relative improvement in this case is still of $14\%$ while it grows to more than $1000\%$ when the amount of data is reduced by 5×5$\times$5 ×.
Figure 4: Accuracy per entity name length for ger-ald versus ger-caption codes. (left): Accuracy averaged per entity name length. (right): Qualitative examples of predictions for long entity names. Code tokens are symbolized between brackets.
Figure 5: ald versus captioning codes. (left): Effect of different code lengths for ger-ald and ger-caption codes. (right): Cumulative distribution function (CDF) of (in green) the position of the least frequent token in the tokenized entity name and of (in pink) the length of tokenized entity name.
Limited model capacity. In Fig. 3 (middle), we see that the model struggles to learn unstructured codes when its capacity is reduced. When considering the small version of our model (114M parameters), the performance with atomic codes is very poor: $0.7$ top-1 accuracy.
Web-scale label-space. In Fig. 3 (right), we vary the number of entities for pretraining. The “positive” entities (see Sec. 4.1) are always included in the pretraining set and the amount of “negative” entities is increased, effectively acting as distractors. First, we see in Fig. 3 (right) that for relatively small-scale label-space ($\leq 100k$), the benefit of having semantic codes versus atomic is small. In this regime we find that the model can memorize all the entities without the need for semantic structure between them. This aligns with the findings of DSI [42]. We evaluate ger further in small label-spaces in Sec. 4.5. However, we see that in million-scale label-space regime, semantic structure becomes important and significantly improves the performance compared to atomic codes: $+26\%$ relative improvement.
Overall, we find that ger-atomic fail to learn unstructured codes when the amount of pretraining data or architecture capacity are reduced, or when the label-space increases to million-scale. Unlike ger-atomic, ger-ald succeed in these scenarios thanks to the semantic structure easing the learning. Next, we analyze how ger-ald improves over another type of semantic codes: ger-caption codes.
4.4.2 ald versus captioning codes
We analyze why unambiguous, language-based and discriminative codes (ger-ald) are more effective for entity recognition than directly decoding the entity name (ger-caption). In Fig. 5 (left), we report the performance of ger-ald and ger-caption when varying the length $L$ of the codes. Fixing a code length $L$ to a caption corresponds to keeping only the first $L^{\text{th}}$ tokens of the entity name. In Fig. 5 (right), we report the cumulative distribution functions (CDF) of (i) the position within the tokenized entity name of the least frequent token among the entire corpus (as described in Sec. 3.2) and (ii) the total number of tokens in the tokenized entity name ($L_{e}$ in the notations of Sec. 3.2)).
Discriminative tokens versus number of tokens. We observe in Fig. 5 (left) that the performance of ger-caption increases drastically from $L=2$ to $L=4$. At the same time, we see in Fig. 5 (right) that for $L=4$, less than half of the entity names are considered in full while more than $80\%$ of the ger-caption codes contain the least frequent token of the entire tokenized name. This hints that what is important for language-based codes is not to describe the full entity name but to include its most discriminative part. We also observe that the performance of captioning increases only moderately from $L=4$ to $L=8$ even though the number of entities considered in full increases drastically from $46.6\%$ to $100\%$. This confirms our intuition that decoding all the entity name tokens does not have a major impact on the performance as long as the most discriminative tokens are decoded. Overall, these observations motivate the ald design of keeping only the most discriminative tokens, which is shown in Fig. 5 to lead to improved performance compared to decoding the full tokenized entity name.
Effect of code length for ger-ald. We see in Fig. 5 (left) that the performance of ger-ald is the best for $L=4$. With smaller code lengths, we need to resort to random tokens a lot to achieve unique codes (see Sec. 3.2), which deters the performance. For example at $L=2$, more than $10\%$ of the entities use a random code token while this percentage decreases to $0.5\%$ at $L=4$. We also see that the performance of ger-ald decreases for code length above $L=4$, which hints that only the few most discriminative tokens are important while additional ones clutter the entity code. Interestingly we also observe in Fig. 5 (left) that when considering all the tokens,ger-ald performance is slightly below that of ger-caption. This might seem surprising since the same amount of information is present in both cases. However we find that when considering all the tokens, it is more difficult for the model to decode tokens ordered by frequencies than tokens ordered syntactically.
Entities with long entity names. In Fig. 4 (left), we report the accuracy per entity name length for both ger-ald and ger-caption finetuned models. We see that the longer the entity name, the more ger-ald improves over captioning. Longer entities tend to have more noise with key information further into the code. We also show in Fig. 4 qualitative examples of entities with long entity names (more in Fig. 12). In the left example, we see that ger-ald use the token combination [col][ob] to represent the semantic concept of colobus monkey species. The last token is used to efficiently differentiate between sub-species of colobus. This compact and discriminative way of encoding the entity allows ger-ald to successfully predict this entity whereas ger-caption fails to generate the entity tokenized name.
Selection strategy HM
Least frequent tokens 14.4
Most frequent tokens 12.3
First tokens 12.0
Random tokens 11.3
Tokens order HM
Least frequent first 14.4
Syntax order 14.4
Random order 13.0
Least frequent last 12.7
Table 3: Ablation study of ger-ald codes. (left) Word tokens selection. (right) Tokens order. All variants use $L=4$. Default is in top rows. Non language-based ger-atomic gets $11.4$ top-1.
Dataset Codes HM
WebLI WebLI caption 1.8
Entity-WebLI (55M) WebLI caption 12.9 (+11.1)
Entity-WebLI (55M) Entity name 14.8 (+1.9)
Entity-WebLI (55M) ALD 17.5 (+2.7)
Figure 6: Entity-based pretraining ablation. (left): Validation OVEN accuracy. (right): Examples of original WebLI captions versus corresponding OVEN entity names.
Figure 7: Pretraining. We vary the size of the pretraining dataset by changing the amount of retrieved examples from WebLI for each OVEN entity (see Sec. 3.3).
4.4.3 Creating codes with ald
Least frequent tokens. In Tab. 3 (left), we validate our choice of selecting the least frequent tokens by evaluating 3 alternatives: random choice, most frequent tokens and first-appearing tokens in the tokenized entity name. We see that these alternative selection strategies hurt the performance significantly. Qualitative examples in Fig. 11 show that the kept tokens are less semantic and discriminative compared to ger-ald strategy of keeping the least frequent tokens. Note that all these variants are at least as good as ger-atomic ($11.4$ top-1) which is not based on language at all.
Decoding order. In Tab. 3 (right), we vary the order of the first $L-1$ tokens in ger-ald codes. Instead of decoding tokens from least to most frequent, we evaluate most to least frequent, syntax order and random order. Note that the selected tokens are the same in all variants, only their order changes. We see that both “least frequent first” and “syntax” orders achieve the best of performance.
4.4.4 Entity-based pretraining
Entity-based pretraining. In Fig. 6, we analyze why our entity-based pretraining improves over the standard captioning pretraining of PaLI or GiT models. First, we see that our method of selecting WebLI data relevant to OVEN entities drastically improves the performance (+11.1 in Fig. 6 (left)). This is because, by design, we select image-text pairs from WebLI that have captions similar to OVEN entity names. Hence, this data is directly relevant for the OVEN entity recognition benchmark. Second, we see that replacing the original WebLI caption with its corresponding entity name from OVEN leads to superior performance (+1.9). We see in the qualitative examples of Fig. 6 (right) that original captions contain a lot of descriptive information not directly relevant to the entity. Lastly, we confirm that using ger-ald codes is better (+2.7) than tokenized entity name.
Dataset size. In Fig. 7, we evaluate the effect of the pretraining dataset size for ger models. We control the dataset size by varying the amount of retrieved examples from WebLI for each of the OVEN entities (see Sec. 3.3). We see in Fig. 7 that ger-ald, ger-caption and ger-atomic benefit greatly from more data and do not seem to have reached saturation yet. As analyzed in Sec. 4.4.1, ger-atomic fails when the amount of pretraining data decreases.
4.5 Link with classification
Method ImageNet-LT WebVision
Classif. MLP 74.3 80.9
ger-atomic $L=2$ 80.8 84.7
ger-ald $L=2$ 80.9 84.8
ger-atomic $L=1$ ($\sim$ Classif. MAP) 81.0 84.8
Previously published numbers
NCR [13] 76.8
CurrNet [10] 79.3
PEL [40] 78.3
MAM [14$\dagger$ 82.3 83.6
Table 4: Evaluation of classification models and ger on small-scale label-spaces. $\dagger$ indicates the use of additional data.
A typical way of tackling visual entity recognition is by training a classifier into the number of entities [35]. This is not a viable solution for web-scale problems such as OVEN where a single fully-connected layer for a $6$M classes has an enormous parameter count of $4.6$B. In this section, we evaluate ger in cases where learning a classification model is a feasible choice (smaller number of classes). Classification can be cast in our ger framework simply by setting $L=1$ and $V=|\mathcal{E}|=$ number of classes (see Sec. 3.1), making it a special case of atomic codes with $L=1$. Since the decoder decodes a single token, it is equivalent to a multi-layer Multihead Attention Pooling (MAP) head [48, 21]. In Tab. 4, we consider two challenging classification datasets: long-tailed ImageNet-LT [24] and noisy Webvision [22]. We evaluate ger-{ald, atomic} and a classification baseline using multi-layer perceptron (MLP) on averaged-pooled patch tokens. Implementation details are in Sec 6.3.
We see in Tab. 4 that using ger-atomic instead of standard MLP improves significantly the performance of the classification model (74.3 versus 81.0 for ImageNet-LT). We also observe that ger-atomic and ger-ald have comparable performance in this relatively small label-space regime (1k classes). As a matter of fact, this achieves state-of-the-art accuracy for both datasets (when no additional external data is used). This shows that ger framework not only excels for large-scale scenarios, but also works well in datasets with smaller number of visual entities, making ger a general framework for visual entity recognition.
5 Conclusion
In this work, we propose a novel generative framework for web-scale visual entity recognition. We represent each entity by a compact, discriminative and semantic code that a generative auto-regressive model learns to decode. In future work, we will explore ways of creating better entity codes by leveraging additional information: either from the Wikipedia page such as the description of the entity and its attached image or also by using external tools.
Acknowledgement.
We thank Xingyi Zhou, Ziniu Hu and Armand Joulin, as well as our teammates for their precious help, support and discussions around this project.
Supplementary Material
Zero-shot Finetuned on OVEN
Method HM Seen Unseen HM Seen Unseen
PaLI-17B [5]  1.8 4.4 1.2 16.0 28.3 11.2
GiT-Large [45]  1.7 4.1 1.2  7.0 17.6 4.3
Table 5: Transferring captioning models to OVEN. We report the harmonic mean (HM) of top-1 accuracy on the seen and unseen test splits for two captioning models: PALI-17B [5] and GiT-Large [45]. Numbers from GiT-Large are run by us. Note that GiT-Large has 42×42$\times$42 × less parameters thank PALI-17B.
6 Implementation Details
We use the Large version of GIT [45] with a pretrained visual encoder and a decoder randomly initialized. The visual encoder is pre-trained with GIT trained for captioning on WebLI dataset [45, 5].
6.1 Entity-based pre-training
We use batch size of 4096, learning rate of $51e-$ 5 for the visual encoder and $1\text{e}^{-4}$ for the decoder, label smoothing of $0.3$ and no weight decay. We use standard inception crop data augmentation for the images. By default and unless specified otherwise, we use code length $L=4$ (see Fig. 5). Note that we only evaluate codes with $L>1$ on OVEN, as the only way to ensure unique codes with $L=1$ is to set $V=|\mathcal{E}|$. This is equivalent to the classification scenario and is not feasible for the million-scale label-space of OVEN. We evaluate $L=1$ in Sec. 4.5 for datasets with a smaller label-space of 1k entities: namely ImageNet-LT [24] and Webvision [22]. Unless specified otherwise our models for the main results (i.e. in Sec. 4.2 and Sec. 4.3) are trained on Entity-WebLI with 55M images ($k=100$) during 600k steps while models for ablations are trained on Entity-WebLI with 27M images ($k=20$) for 200k steps.
Preventing data leakage. Webli is already deduplicated against the train, val, and test splits of 68 common vision/vision-language datasets (see PaLI paper [5]). To be sure, in our paper, we further removed pretraining images with a cosine similarity (with CLIP-L/14 visual features) above 0.95 with any of the OVEN images. We chose a 0.95 conservative threshold by looking at some examples: similarity 0.95 corresponds to conceptually similar images but clearly not duplicates (see Fig 8).
Figure 8: Filtering out pretraining data too similar to OVEN test/val.
Algorithm 1 ger-ald codes.
Data: Code length L, Text tokenizer $\Phi(.)$, Entities $\mathcal{E}$
Result: $\mathcal{C}=\{c_{e}\}_{e\in\mathcal{E}}$
for $v\in[1,V]$ do
6.2 Finetuning on OVEN train set
We finetune models on OVEN training set for 30,000 steps with a batch size of 256 and a learning rate of $1\text{e}^{-7}$. Label smoothing is set at $0.1$. Note that the finetuning schedule is relatively short (30,000 steps) because we observe that long finetuning (or equivalently, using a large learning rate) causes the model to forget about the unseen categories.
6.3 Training on ImageNet-LT and Webvision
We train the model on ImageNet-LT and Webvision datasets with the batch size of 512 and a learning rate of $1\text{e}^{-4}$ for both encoder and decoder. We do not use any label smoothing but apply a dropout of $0.1$. We use $L=2$ for these experiments with ger-ald because unlike very large label-space this is enough not to resort to random tokens when ensuring that codes are unambious.
Figure 9: Zero-shot versus finetuned captioning models predictions. We qualitatively compare the predictions of the captioning GiT-Large model when evaluated on OVEN in a zero-shot manner or after finetuning on OVEN train set.
6.4 Implementation details about the baselines
Dual encoder with CLIP-L/14.
We use a learning rate of $3\text{e}^{-7}$, a batch size of $4096$ and we train for $200,000$ steps since training for longer deteriorates the performance. During finetuning on OVEN training set, we find it important to still include some pretraining data: we randomly sample, with a probability of $90\%$, elements from the pretraining dataset. Otherwise, when finetuning solely on OVEN, the model becomes too specialized for the seen categories and is not capable of discriminating between all the negative entities. Alternatives could be to freeze some layers of the network during finetuning to prevent catastrophic forgetting.
ger-atomic.
We benchmark different values for the choice of $L$: {$2$, $4$, $8$} and $V$: {$512$, $4096$, $32768$} when using atomic codes. Our default is to use $L=2$ and $V=4096$. Note that this corresponds to more than $16$M possible different codes and we use only a subset of $6$M of those unique codes.
ger-hkc.
For HKC, we first represent each Wikipedia entity with a text embedding using the sentence-t5 [28] encoder. We have experimented with different ways of creating such embeddings, such as creating text embeddings from the Wikipedia titles, Wikipedia article summaries, and Wikipedia title and article summaries combined together. We have observed that Wikipedia title and article summaries combined together produces the best embeddings. We have tried different values of $k$: {$10$, $100$, $1000$, $4096$, $8142$}, and found that $k=4096$ achieves the best performance in the validation set.
7 More Experimental Results
7.1 Failure case analyses
We show in Fig. 10 that our method works well across different entity types. In many cases, our codes for animals (e.g. ‘Glaucous winged gull’), persons (e.g. ‘List of celebrities who own wineries and vineyards’) or organizations (e.g. Gladney Center for Adoption’) are interpretable, as shown in the qualitative examples in Fig 11 and Fig 12. We see failure cases when semantics is difficult to infer from entity name alone. This is often the case for scientific denomination of species as show in Fig. 10. In future work, using external tools or Wikipedia page content could improve results in such cases.
Figure 10: (left): Accuracy per sub-task in OVEN. (right): A failure case example in iNaturalist (‘inat’) sub-task.
7.2 Zero-shot OVEN with captioning models
Quantitative evaluation.
In Table 5, we transfer two captioning models, namely PALI-17B and GiT-Large [45], both pre-trained on WebLI [5] to the OVEN task. We observe in Table 5 that these models transfer poorly in a zero-shot manner. This can be explained by the major discrepancy between the pre-training captioning task (i.e. describing an image with a caption) and the target entity recognition task.
Visual examples.
We show some visual examples of predictions from the validation test between the zero-shot and finetuned GiT-Large models in Figure 9 where we clearly see the difference of output between the zero-shot and finetuned GiT-Large models. In the left column, we show two examples where both zero-shot and finetuned models fail. However, even though the finetuned model fails to find the correct category of castle or plant, it still tries to output a fine-grained category of castle or plant. This is not the case of the zero-shot model which gives a generic description of the entity, for example “The castle in the middle ages.”. In the middle column, we show examples where zero-shot fails but the finetuned model finds the correct category. Finally, in the last column we show cases where the zero-shot model succeeds, but even when it does we observe that the generated caption is cluttered (for example with “a photo of a picture”) while the finetuned model directly outputs the entity name.
Overall, the observation that models pre-trained from WebLi captions do not generalize well to OVEN entity recognition motivated us to create our entity-based pre-training described in Section 3.3.
Figure 11: Token selection strategies in ger-ald. We qualitatively compare different alternative token selection strategies for ger-ald: most frequent token or random token selection. We use $L=2$ for this qualitative evaluation since this is easier to visually interpret that $L=4$, however the trends are consistent. Quantitative evaluation is in the Table 3 of the main paper.
7.3 Entities with long names
In Figure 12, we show more visual examples of ger-ald and ger-caption predictions for entities with long names.
Figure 12: Qualitative study of ger-ald versus ger-caption. Visual examples of predictions for long entity name from 9 to 16 tokens. For these visualizations with ger-ald, we use SentencePiece tokenizer [20] and $L=2$ in this evaluation since this leads to more visually interpretable codes. Tokens are symbolized between brackets. We report the top-3 predictions for ger-ald and for ger-caption codes, and color in green the correct predictions. We observe that ger-ald codes are easier to predict as they contain less clutter than ger-caption codes. Interestingly, we see that with ger-ald the top-3 predictions usually share a common token which re-group different semantically close entities.
Pretraining size (M) 10.6 14.7 26.6 40.3 54.9
ger-atomic 0.9 6.8 11.4 13.8 15.3
ger-ald 10.2 11.7 14.4 16.1 17.5
Relative $\Delta$ (%) 1029 71 26 17 14
Architecture size (M) 114 179 397
ger-atomic 0.7 5.4 11.4
ger-ald 5.6 9.5 14.4
Relative $\Delta$ (%) 648 77 26
# entities (M) 0.02 0.03 0.12 1.00 6.08
ger-atomic 34.6 34.0 29.7 21.5 11.4
ger-ald 33.5 33.5 30.4 23.6 14.4
Relative $\Delta$ (%) -3.1 -1.6 2.2 9.6 26.4
Table 6: Semantically-structured (ger-ald) versus unstructured (ger-atomic) codes. We report the numbers corresponding to Figure 3 of the main paper. The pretraining dataset sizes of 10.6M, 14.7M, 26.6M, 40.3M and 54.9M correspond respectively to setting $k$ to 2, 5, 20, 50 and 100. The architecture sizes with 114M, 179M and 397M parameters correspond respectively to variant Small, Base and Large of the model. The label space sizes with 20549, 30549, 120549, 1000549 and 6084491 different entities correspond respectively to having 0, 10k, 100k, 1M and 6M entities acting as distractors.
7.4 Different token selection strategies
In Figure 11, we show visual examples of codes generated with alternatives of selecting the least frequent token in ger-ald. We compare with selecting instead the most frequent token and with selecting a random token in the entity name. Quantitative evaluation is in Table 3 of the main paper. In Figure 11, we observe that codes generated with least frequent token strategy are the most semantically structured. Indeed, in this case the entities “Adoption by celebrities” and “List of celebrities who own wineries and vineyards.” share a common token (the token corresponding to “celebrities”) while there is no intersection of token between those two entities for the most frequent or the random strategies. We observe the same effect across several group of entities that we intuitively expect to have shared tokens, for example with “smoking and pregnancy”, “teenage pregnancy in Australia” and ”Immunization during pregrancy”, or with “Denall National Park and Preserve” and “Bishop Ranch Regional Preserve”.
7.5 Numbers corresponding to Fig. 3 main paper
We report in Table 6 the numbers corresponding to the experiments shown in Figure 3 of the main paper.